### v1

# 1st Xing Fang College of Computer and Information Engineering

Abstract
The human immune response depends on the binding of T-cell receptors (TCRs) to antigens (pTCR), which elicits the T cells to eliminate viruses, tumor cells, and other pathogens. The ability of human immunity system responding to unknown viruses and bacteria stems from the TCR diversity. However, this vast diversity poses challenges on the TCR-antigen binding prediction methods. In this study, we propose a Masked Language Model (MLM), referred to as tcrLM, to overcome limitations in model generalization. Specifically, we randomly masked sequence segments and train tcrLM to infer the masked segment, thereby extract expressive feature from TCR sequences. Meanwhile, we introduced virtual adversarial training techniques to enhance the model’s robustness. We built the largest TCR CDR3 sequence dataset to date (comprising 2,277,773,840 residuals), and pre-trained tcrLM on this dataset. Our extensive experimental results demonstrate that tcrLM achieved AUC values of 0.937 and 0.933 on independent test sets and external validation sets, respectively, which remarkably outperformed four previously published prediction methods. On a large-scale COVID-19 pTCR binding test set, our method outperforms the current state-of-the-art method by at least 8%, highlighting the generalizability of our method. Furthermore, we validated that our approach effectively predicts immunotherapy response and clinical outcomes on a clinical cohorts. These findings clearly indicate that tcrLM exhibits significant potential in predicting antigenic immunogenicity.
The source code and datasets used in this study are available at: https://github.com/hliulab/tcrLM.
Index Terms:
Immune Response, Large Language Model, Transformer, Virtual Adversarial Training, T Cell Receptor, Tumor Antigen
I Introduction
Antigen recognition is a prerequisite for triggering the T cell-mediated anti-cancer immune response. T cells interact via a dimeric surface protein, the T-cell receptor (TCR), with an antigen presented on a major histocompatibility complex (MHC) located on the surface of antigen-presenting cells (APC). Their interaction triggers the activation of T-cells, leading to their clonal expansion and differentiation into effector T-cells, such as cytotoxic T lymphocytes (CTLs). These CTLs can migrate to the tumor microenvironment and directly kill tumor cells through secreting toxic molecules, like perforin and granzyme, or indirectly promote tumor cell death by modulating the immune environment. In addition, a portion of activated T-cells further differentiate into memory T-cells, which survive long-term in the body, providing the host with long-term immune memory and a rapid response capability against tumors.
The T-cell receptor (TCR) comprises $\alpha$ and $\beta$ chains that are generated through genetic recombination, leading to an expansive TCR repertoire. Previous studies have documented that humans have the potential to generate approximately 1015 to 1020 distinct TCR sequences. This diversity predominantly manifests in the complementarity determining region 3 (CDR3) [1], which engages directly with the peptide-MHC complex, thereby dictating the binding specificity of the TCR [2, 3]. The antigens recognized by TCRs are peptide segments composed of 8-14 amino acids. Essentially, both the TCR CDR3 and antigens can be expressed as a sequence of amino acids, exhibiting striking similarity to human natural language. Consequently, we can draw inspiration from methods in the field of in Natural Language Processing, and utilize advanced language models to learn and analyze these biological sequences.
This marks the first application of large language model technology in the field of immunology, promising unprecedented breakthroughs, a deeper understanding of TCR complexity, and potentially opening a new chapter in immunotherapy.
II Related Work
II-A Predicting pTCR binding specificity
Quite a few computational methods have been developed to predict the binding specificity of T-cell receptors (TCRs) to peptide-MHC complexes (pMHCs). These methods roughly fall into three main categories: 1) Clustering-based methods measure the similarities between TCRs and try to understand underlying patterns in binding to antigens. Representative methods include TCRdist [4], DeepTCR [5], GIANA [6], iSMART [7], GLIPH [8], and ELATE [9] models. 2) Peptide-specific models focus on predicting the binding of specific peptide segments to TCRs, including TCRGP [10], TCRex [11], and NetTCR-2 [12]. 3) Generic binding prediction models are not limited to specific peptides but require models trained on known TCR bindings, including PanPep [13], pMTnet [14], DLpTCR [15], ERGO2 [16], and TITAN [17]. Although these methods have demonstrated promising accuracy in specific scenarios, they have limitations in generalizing to unseen peptides, which is crucial for identifying the binding specificity of neoantigens or exogenous antigens. Therefore, accurately identifying the pTCR bindings remains a challenging task.
(a) TCR Frequency
(b) Antigen Frequency
Figure 1: Statistical frequency distributions of TCR and antigen sequences with respect to length
II-B Protein language model
Large language models have gained increasing attention in the field of protein modeling in recent years, and yield to new insights and capabilities for understanding the structures and functions of proteins. Most protein language models are built upon the Transformer encoder, which can encode protein sequences or structures into fixed-length latent representations. The representations have been proven to boost the performance of downstream tasks related to proteins, such as the prediction of structure, functions, and protein-drug interactions. The masked language models (MLM), such as ESM1b [18], ESM-1v [19], ProtFlash [20], and ProtTrans [21], aim to reconstruct masked tokens based on surrounding sequences. Their success in protein tasks has inspired us to make advantage of MLM to accommodate the diversity of TCR sequences.
(a) Masked language model
(b) pTCR prediction network
Figure 2: Illustrative diagram of tcrLM model and its encoder architecture, as well as the network for pTCR binding prediction
III MATERIALS AND METHODS
III-A Data source and preprocessing
For model pre-training, we set about to establish a largest TCR CDR3 sequences to date. For this purpose, we collected 113,888,692 distinct TCR CDR3 sequences from more than ten databases and publications. With over 100 million sequences, our model is able to capture underlying feature from TCR sequences. The frequency distribution of TCR CDR3 sequences is illustrated in Fig. 1(a), which reveals that most CDR3 sequences range between 10 to 20 amino acids in length.
For fine-tuning, we build a benchmark dataset of pTCR bindings by collecting a substantial amount of pTCR binding data from various databases. We consider both the $\alpha$ and $\beta$ chains of TCR and treat them as independent CDR3 sequences, as previous studies have demonstrated both of them play critical role in antigen recognition. The pTCR dataset consists of 109,554 bindings, covering 1,377 unique antigens and 104,623 unique TCR CDR3 sequences. To our best knowledge, this the largest pTCR binding dataset to date.
The frequency distribution of antigen sequences is shown in Fig. 1(b), indicating that most antigens 9mer in length. Next, we randomly mismatched TCR and peptide sequences to generate the same number of negative pTCR samples. During the fine-tuning stage, we randomly select 10% of the pTCR binding dataset as an independent test set, while the remaining 90% is used as the training set. The training set encompassed 881 distinct antigens and 93,727 unique CDR3 sequences, while the test set included 496 distinct antigens and 10,896 unique CDR3 sequences. The five-fold cross-validation is used to optimize the model hyperparameters.
III-B tcrLM model
The BERT model  [22] pre-trains deep bidirectional representations through the Masked Language Model (MLM) task, which allows to freely capture information from all tokens in the context, whether they appear before or after the current word. Therefore, we apply the BERT-based language model, named tcrLM, to utilize Transformer encoder and its training methodology for the masked language model to effectively process and parse TCR sequences.
As illustrated in Fig. 2(a), each TCR CDR3 sequence is padded to a fixed length of 20, since all CDR3 sequences do not exceed 20. We randomly generate a mask of length between 3 to 5 to cover a continuous segment in the TCR sequence. The masked sequence passes through an embedding layer that converts each token into a high-dimensional embedding. The embeddings are then taken as input to the encoder and ultimately mapped into a 20x512 matrix, aiming to capture the underlying biological features within the TCR sequences. Next, the matrix is flattened into a 2176-dimensional vector, which subsequently passes through three fully connected layers with node counts of 256, 64, and 420 (20x21), activated by the ReLU function. As a result, the vector is reshaped into a 20x21 matrix that represents the amino acid distribution at each position of a TCR sequence. We introduced the rotary positional embedding (RoPE)[23] into tcrLM. This method encodes absolute positions using rotation matrices and integrates explicit relative positional dependencies into self-attention calculations. RoPE offers several advantages, such as flexibility in sequence length, decaying dependencies between tokens as their relative distance increases, and the capability to enhance linear self-attention with relative positional encoding. Furthermore, we incorporated the mixed chunk attention (MCA) [24] to combine the advantages of both local attention and linear attention mechanisms.
The language model is trained to recover the masked segment using the surrounding sequences on the collected more than 100 million TCR sequences. In our practice, we applied different learning rates to distinct components of the model. For the encoder, the learning rate was set to 1e-6, while for the subsequent fully connected layers, it was adjusted to 1e-4. The optimizer was set to Adam. The pre-training was conducted on a CentOS Linux 8.2.2004 (Core) system, equipped with an Intel(R) Xeon(R) Silver 4210R CPU running at 2.40GHz, along with a GeForce RTX 4090 GPU and 128GB of RAM. The model was implemented using the PyTorch 2.2.1 framework, and accelerated by the DeepSpeed 0.13.5 library. The entire training process spanned four days, resulting in a reduction of the perplexity from 10.4 to 6.4.
III-C Predicting pTCR binding
The pre-trained encoder is used in the prediction network for pTCR binding. As shown in Fig. 2(b), the antigen sequences and TCR sequences are padded to the maximum length of 15 and 20, respectively. After passing through an embedding layer, these two sequences are taken as input of the pre-trained encoder, and transformed into matrices of dimensions 15x512 and 20x512. Subsequently, these two matrices are concatenated along the feature dimension to form a 35x512 matrix, which is then flattened into a 17,920-dimensional vector. This vector passes through a fully-connected layer and is projected to a 2-dimensional vector, followed by softmax layer to obtain the final predicted probabilities. During the training process, the pre-trained encoder is frozen. For pTCR binding prediction task, we use the cross-entropy as the loss function:
$\mathcal{L}_{cls}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(p_{i})+(1-y_{i})\cdot\log(1-p_{i})\right]$ (1)
in which $y_{i}$ and $p_{i}$ represent the actual and predicted pTCR binding, and $N$ is the total number of training samples.
(a) ROC curves
(b) Performance comparison
(c) PPV
Figure 3: Performance evaluation of tcrLM and four comparative methods on independent test set
III-D Virtual adversarial training
Given the vast diversity of the TCR repertoire, the training data currently available is still limited and even biased. This poses a tough challenge on the development of robust prediction models for pTCR binding. To address this challenge, we employs virtual adversarial training to enhance the model’s generalizability. Since protein sequence are discrete, we introduce perturbations in the embedding layer of the sequences [25]. For this purpose, we create adversarial examples designed to maximize the loss function by applying adversarial perturbations to the sequence embeddings, namely, the adversaries are generated in the direction of gradient ascend and constrained by L2 norm. Formally, we define the adversarial loss as below:
$L_{\mathrm{vadv}}(x,\theta)$ $=D\left[p(y|x,\hat{\theta}),p(y|x+r_{\mathrm{vadv}},\theta)\right]),$ (2)
[…]
$D$ represents the function that measures the divergence between two distributions, $p(y|x)$ denotes the probability of the model predicting label $y$ given input $x$, $r_{\mathrm{vadv}}$ is a virtual adversarial perturbation regarding the input sample $x$. This perturbation strives to maximize the divergence between $p(y|x_{*},\hat{\theta})$ and $p(y|x+r)$ by choosing the direction of gradient ascent.
Virtual adversarial training requires the model to minimize not only the risk on actual observed data but also the risk arising from adversarial loss. This approach helps reduce the model’s sensitivity to slight input variations, thereby enhancing its generalizability. By incorporating virtual adversarial training, we aim to build a more robust and reliable prediction model, capable of handling the vast diversity of the TCR repertoire even with limited and potentially biased training data.
IV Results
IV-A Evaluation on independent test set
We first assessed the performance of tcrLM on the 10% hold-out independent test set. To benchmark the performance in predicting pTCR binding specificity, we conducted a comparative analysis with four currently state-of-the-art methods: PanPep [13], ERGO2 [16], pMTnet  [14], and DLpTCR [15]. For the former three methods, we executed their executable codes using their recommended parameters on the same workstation. For DLpTCR, we utilized its web server to make predictions on the test data.
We found that tcrLM significantly outperforms all other methods, as shown in Fig. 3(a-c). Specifically, tcrLM achieved an AUROC of 0.937 and an AUPR of 0.933, highlighting its exceptional predictive capability in determining pTCR binding specificity. Among the compared methods, only ERGO2 demonstrated moderate performance, with an AUROC of 0.704 and an AUPR of 0.747. The remaining methods performed close to random guessing, indicating their limited performance in predicting pTCR binding. These previous methods may showed promising performance on small datasets, but their performance declined significantly when tested on large-scale datasets. This suggests that they have weak generalizability and struggle to adapt to real-world, large-scale data scenarios.
To further examine our model’s ability to prioritize pTCR bindings, we calculated the positive predictive value (PPV) for the top-ranked predicted pTCR samples on two different datasets. Specifically, we evaluated the PPV for the top 100, top 1000, and top 5000 predictions. As shown in Fig. 3(d), tcrLM achieved PPV values of 99%, 98.7%, and 96.74% for the top 100, top 1000, and top 5000 predictions, respectively. In comparison, other methods demonstrated inferior prioritization capacity compared to tcrLM.
(a) Performance on external set
(b) Performance on COVID-19 set
(c) PPV
Figure 4: Performance evaluation of tcrLM and four comparative methods on external and COVID-19 test set
IV-B Evaluation on external set
To further evaluate the performance of our model, we constructed an external test set collected from tens of publications. The external set included 63,324 pTCR binding samples across 998 unique peptides and 53,439 CDR3 sequences, which allowed us to assess the capacity of our method in real-world scenarios. As shown in Fig. 4(a), tcrLM achieved AUROC and AUPR values exceeding 0.9, and significantly outperformed the second-best method, ERGO2, which had AUROC and AUPR values of only about 0.7. This observation reflected the limitations of previously published methods, while in turn validated the superior performance of tcrLM in predicting pTCR binding specificity between unseen peptides and TCR sequences.
IV-C Performance evaluation on COVID-19 dataset
To validate the generalizability of tcrLM, we tested its capacity to predict binding between virus-derived antigens and TCRs. We collected a total of 520,000 binding samples between COVID-19 virus antigens and human TCRs from the ImmuneCODE database [26]. For class balance, we generated equal number of negative samples through random shuffle. As a result, we created a million-scale COVID-19 test set, which is the largest pTCR binding test set to date.
For objective performance evaluation, we again compared tcrLM with previously published methods, including PanPep, ERGO2, DLpTCR, and pMTnet [14]. As shown in Fig.4(b), tcrLM achieved AUROC and AUPR values of 0.595 and 0.602, respectively, while the comparative methods obtained AUROC and AUPR values only slightly above 0.5, which is nearly close to random guessing. Furthermore, we calculated the positive predictive value (PPV) of the top 100, top 1000, and top 5000 predictions made by each method. Our model consistently achieved 93% PPV value, significantly better than all competitors, whose PPV values remained below 65% (Fig.4c). Overall, the significant advantage over previous methods strongly validates the robust generalizability of tcrLM and highlights its potential in facilitating effective immune-based therapies and vaccine design targeting the COVID-19 virus.
IV-D tcrLM predicts immunotherapy outcomes
To further explore the predictive capacity of tcrLM, we conducted an exploratory analysis of pTCR in a cohort of patients with advanced melanoma. This cohort consists of 29 patients who underwent treatment by immune checkpoint inhibitors. Based on the TCR-seq and genomic sequencing data, we took the missense mutations as anchors, and generated all possible 9-mer peptides covering the anchors. After extracting CDR3 sequences from the TCR-seq data, we created all conceivable peptide-CDR3 pairs for each patient, yielding a total of 81,851,486 pTCR candidates.
We scored these samples using tcrLM and selected the top 25,000 highest-scored pairs for each patient. Subsequently, we categorized the patients into four categories according to RECIST criteria: Complete Remission (CR), Partial Remission (PR), Stable Disease (SD), and Progressive Disease (PD) groups. As shown in Fig. 5(a), the boxplots of predicted scores of four groups differed significantly from each other. The analysis of variance verified the statistical differences among the groups (F-test, p-value¡0.01). By stratifying the patients into benefit, non-benefit, and long-term survival groups, we observed that patients in the long-term survival group had higher predicted scores compared to the other groups (Fig. 5b). This findings validated the correlation between higher pTCR predicted scores and better immunotherapy outcomes, offering a new perspective and a possible clinical indicator for tumor immunotherapy.
(a) Boxplots of RECIST grouping
(b) Violin plots of long-term grouping
Figure 5: Combined Box Plot and Violin Plot
IV-E Ablation experiments
To verify the effectiveness of the pre-trained encoder, we conducted ablation experiments to evaluate its impact on the performance in predicting pTCR binding. In particular, we removed the pre-trained encoder as the antigen sequence encoder, the TCR sequence encoder, or both. The performance comparison results of the three ablated models were shown in Table  I. We found that removal of the pre-trained encoder from the model always led to performance decline. In particular, when both the antigen and TCR encoders were removed simultaneously, the AUC value dropped to 0.85 AUC value, demonstrating the significance of the pre-trained encoder.
We also tested the impact virtual adversarial training, and found that the ablated model without virtual adversarial led to 2% performance reduction.
TABLE I: Performance evaluation of ablated models
Ablated model AUROC Accuracy MCC F1 AUPR
tcrLM 0.9404 0.8612 0.7234 0.8574 0.9331
Without TCR encoder 0.9213 0.8474 0.6998 0.8554 0.9213
Without antigen encoder 0.8617 0.7905 0.5812 0.7880 0.8551
Without TCR&antigen Encoder 0.8522 0.7737 0.5478 0.7702 0.8765
Without Virtual Adversarial 0.9266 0.8390 0.7082 0.8508 0.9180
V Conclusion
This study proposed a BERT-based large language model, named tcrLM, to enhance the accuracy of predicting pTCR binding specificity by utilizing the strong capability in feature extraction of large language models. We have pre-trained the model on more than 100 million TCR sequence, and verified its performance on downstream tasks. Our experimental results demonstrated that tcrLM exhibited superior performance compared to the current four state-of-the-art models, on both independent and external test sets. Moreover, in the real-world scenario of COVID-19 and patient-derived datasets, tcrLM still yielded promising performance, highlighting its potential applicability in vaccine design and immunotherapy treatment.